OpenStudy (anonymous):

At what points will the line y = x intersect the unit circle x2 + y2 = 1?

OpenStudy (anonymous):

since \(y=x\) you can solve \[x^2+x^2=1\] for \(x\)

OpenStudy (anonymous):

takes only a couple of steps, but be careful \[2x^2=1\\x^2=\frac{1}{2}\\x=\pm\frac{1}{\sqrt2}\]

OpenStudy (anonymous):